OpenStudy (anonymous):

Which of the following points lies on the circle whose center is at the origin and whose radius is 10? (-6, 4) (√(10)√(10)) (6, -8)

OpenStudy (anonymous):

Draw a right triangle inside the circle with the radius of the circle as the hypotenuse. You can use the pythagorean theorem, which says that x^2 + y^2 = c^2 (with c being the hypotenuse of the triangle). So you will have a^2 + b^2 = 10^2= 100 you just have to find values for a and b that work. So try, does (-6)^2 +(4)^2 = 100? 36+ 16= 52. NO Then does √(10)^2 + √(10)^2 = 100? 10+10= 20. NO What about 6^2 + (-8)^2= 100? 36+64=100 YES. The correct answer is this one